OpenStudy (anonymous):

Your company needs to temporarily hire a programmer to work on a project. Two proposed payment schemes for this work are as follows: • (1) A flat fee of $1,000, plus $20 per hour or (2) $25 per hour. o For each of the two plans, show an expression that can be used to compute the amount of pay for that plan. The variable should be the number of hours worked. o Set up and solve an inequality that would enable your company to determine possible job lengths (in hours) for which the person is paid less according to plan 1 than for plan 2.

hero (hero):

y = 1000 + 20x y = 25x That's the first part.

hero (hero):

25x < y < 1000 + 20x That might be a formula for plan two

OpenStudy (anonymous):

ok thank you!

hero (hero):

Actually, I think you have to solve each for x and then write the inequality.

OpenStudy (anonymous):

i think you are correct

hero (hero):

But we can actually do it ourselves and find the exact number of hours it would take. Let's first write each in terms of x...

OpenStudy (anonymous):

i came up with plan one will give him 1800 but im still working on plan 2

OpenStudy (anonymous):

ok and i see he only gets 1000 from plan 2

hero (hero):

x = (y - 1000)/20 x = y/25 y/25 < x < (y - 1000)/20 x = x 25(y - 1000) = 20(y) 25y - 25000 = 20y 5y = 25000 y = 25000/5 y = 5000

OpenStudy (anonymous):

ahh but without that 1k bonus plan 1 will get him less money in th elong run

OpenStudy (anonymous):

oh! lol i didnt calculate by the year lol i only did it for the first paycheck

hero (hero):

x = 4000/20 x = 200

hero (hero):

So 200 hours is the point at which everything switches as far as the plans go. If you go over 200 hours, then the plans will likely switch.

OpenStudy (anonymous):

ahh i see! thanks hero! it makes sense now

hero (hero):

Since he's only temporary, he might pick the first plan. He might not get to 200 hours.

hero (hero):

It will take him 25 days to get to 200 hours.

hero (hero):

So, if he plans to work there for over a month, he should pick plan 2. For less than a month, stick with the first plan.

hero (hero):

There's your complete solution.

hero (hero):

You should keep these steps in mind. You may have to make the same decisions one day.

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

lmao i know right XD no bonuses for me lol

hero (hero):

keep your hopes up

hero (hero):

Romie, are you still around?

hero (hero):

Definitely good questions